# OnlineSI: Taming Large Language Model for Online 3D Understanding and Grounding

### v1

@@ Abstract
In recent years, researchers have increasingly been interested in how to enable Multimodal Large Language Models (MLLM) to possess spatial understanding and reasoning capabilities. However, most existing methods overlook the importance of the ability to continuously work in an ever-changing world, and lack the possibility of deployment on embodied systems in real-world environments. In this work, we introduce OnlineSI, a framework that can continuously improve its spatial understanding of its surroundings given a video stream. Our core idea is to maintain a finite spatial memory to retain past observations, ensuring the computation required for each inference does not increase as the input accumulates. We further integrate 3D point cloud information with semantic information, helping MLLM to better locate and identify objects in the scene. To evaluate our method, we introduce the Fuzzy $F_{1}$-Score to mitigate ambiguity, and test our method on two representative datasets. Experiments demonstrate the effectiveness of our method, paving the way towards real-world embodied systems. The project page is available at https://onlinesi.github.io/.
@@ 1 Introduction
However, equipping Multimodal Large Language Models (MLLM) with this form of online spatial intelligence remains challenging. First, most existing methods face prohibitive computational scaling as input stream grows. The architectures that repeatedly process all past observations with full attention layers quickly exhaust the model context and computational budget due to spatial and contextual redundancy. Second, many models still fail in understanding spatial relationships and reasoning in space, a limitation that is further exacerbated by the scarcity of 3D-grounded training data. While recent works, e.g., Hu et al. [20], have explored memory management for embodied planning, their solutions have two key limitations. As task execution proceeds, the memory bank infinitely grows, leading to computational and storage bottlenecks. Furthermore, their model perceives 3D space at a coarse-grained level, rendering them unsuitable for guiding fine-grained operations such as precise object manipulation. These difficulties motivate our core research question: Can multimodal large language models perceive and understand the 3D world in an online fashion?
In this paper, we propose OnlineSI, a framework designed to understand 3D scenes in an online fashion and make detections on the fly from streaming video. Our key insight is to maintain a finite, explicit spatial memory, and leverage the multimodal model to reason over this compact memory representation for precise, object-level predictions. This design ensures a fixed upper bound for the computational budget of each inference step, effectively decoupling inference cost from the stream length. At the same time, the explicit spatial memory structure also facilitates the model’s grasp of fine-grained spatial concepts. Furthermore, we introduce a novel fusion technique that tightly integrates 3D point cloud data with semantic information, providing the MLLM with a rich, object-level understanding of the scene.
@@
Evaluating performance on an arbitrary video stream introduces a significant ambiguity: given the inherently incomplete observation, should a partially viewed object be counted as a detection? To address this evaluation challenge, we propose the Fuzzy $F_{1}$-Score, a modification of the standard $F_{1}$-Score designed for this online, partial observation setting. We conduct extensive experiments on two representative datasets, ScanNet [10] and ScanNet++ [58]. Experimental results show that our method significantly outperforms existing baselines, demonstrate its effectiveness in the online setting.
@@
  • We introduce OnlineSI, a novel framework for online 3D scene understanding and grounding that maintains bounded inference costs, enabling it to process video streams incrementally.
@@ 2 Related Work
3D Object Detection aims to predict category labels and oriented 3D bounding boxes for each instance in the scene. Previous work can be divided into two main genres: methods that operate directly on 3D point clouds and those that use single-view 2D images. In the point-cloud-based genre, research has focused on enhancing the model’s understanding by enriching the 3D representation. A prominent strategy is to leverage knowledge from other modalities. For instance, [36, 37, 4, 5] tap into knowledge priors from 2D images to achieve open-vocabulary 3D detection. Similarly, [40] employs a Large Language Model (LLM) to better process global-local information. Other approaches improve the model architecture or training strategy. For example, [46] uses a vertex relative positional encoding method to help the model focus on points near objects, while [26] improves the system’s generalization through multi-dataset training. In the single-view-image-based genre, [3] pioneers unified monocular 3D object detection by training on the extensive Omni3D dataset. Following this, [31] introduces the first successful BEV-based monocular 3D object detector. A central theme in this area is addressing the scarcity of 3D training data by leveraging abundant 2D data. To this end, [21] uses open-vocabulary 2D models and pseudo-LiDAR to automatically label 3D objects. [57] leverages 2D results from Grounding DINO [34] with a 3D head to predict 3D bounding boxes. Both [24] and [60] also address data scarcity by generating pseudo-3D data from 2D annotations or leveraging pre-trained 2D model knowledge. In our work, we emphasize on online detection from a video stream. A work with similar topic is [63], which unifies grounding and exploration for embodied navigation. Compared to their work, we train a large language model to directly understand 3D point clouds and output 3D bounding boxes with rotations and class labels.
@@
Multimodal Large Language Models aim to understand, reason, and generate from multimodal information. Early methods, such as [42], unifies images and texts by learning a joint image-text representation. More recently, with the remarkable success of Large Language Model and scaling laws, innumerable works have been devoted to unify vision and language within a single large transformer-based model [29, 1, 28, 33, 52, 12, 2, 8, 13]. Meanwhile, others have attempted to unify multiple modalities, enabling models to understand diverse information sources [23, 25, 53]. With the rapid development of autonomous driving and embodied AI, 3D understanding has recently become a significant focus for MLLM researchers. Some works aim to enable models to perceive spatial relationships from images by training on datasets containing 3D annotations [7, 9, 38, 54], while others seek to adapt models to 3D representations for spatial reasoning [16, 41, 55, 62, 14, 20, 50, 61, 22]. However, most of the previous methods suffer from computational overload as the number of input images increases, making them unsuitable for deployment in embodied systems that need to learn at test time. Although some work has explored how to maintain a spatial memory for scene understanding [56, 64], they do not focus on continuously improving understanding and online responding. In this sense, the work most similar to ours is [20]. However, they still have problems such as memory bank growing and coarse-level understanding. In our work, we update an explicit spatial memory based on input frames and use this memory as input to the MLLM, which guarantees a bounded inference time and fine-grained understanding.
@@
2D Object Detection is one of the most fundamental and challenging problems in computer vision. In earlier years, researchers typically use carefully designed, handcrafted features for detection [48, 11, 15]. Then with the rise of deep convolutional neural networks (CNN), many works introduced deep learning to this field, and they can be divided into two groups. The first one follows a “coarse-to-fine” process, proposing coarse regions and then focusing on them [17, 19, 44]. The second group, in contrast, retrieves all objects in one-step inference [43, 35, 32, 27, 6]. Recently, significant research has focused on open-vocabulary object detection, often achieved through large-scale pre-training [59, 18, 30, 34]. In our work, we use an off-the-shelf open-vocabulary detector to obtain pixel-level semantic labels. This semantic information is then injected into the point cloud to enhance its representation.
@@ 3 Preliminaries: SpatialLM
We leverage SpatialLM’s [39] ability of understanding point clouds to achieve online spatial intelligence. SpatialLM is a multimodal large language model designed to understand 3D indoor environment and generate structured scene descriptions. The model employs Sonata [51] as its point cloud encoder and finetunes Llama-3.2-1B-Instruct [47] on a synthetic dataset consisting of 12,328 distinct scenes.
@@ 4.1 Problem Formulation
Given a potentially infinite sequence of images $\{\mathbf{I}_{t}\in\mathbb{R}^{H\times W\times 3}\}_{t=1}^{N}$, for each current image $\mathbf{I}_{t}$, our goal is to predict the 3D bounding boxes and class labels for objects that have appeared in the scene, represented in a unified coordinate frame. As the system must be capable of long-term learning from the environment, we impose two constraints: the inference computational budget per frame must be capped, and the large language model’s input must not exceed its context limit. Therefore, we need to dynamically maintain a memory $\mathbf{M}_{t}$ of past observations, subject to an upper bound on its size.
@@ 4.2 Spatial Memory Management
where $\mathbf{P}_{t-1}$, $\mathbf{X}^{\prime}_{t}$, $\mathbf{S}_{t-1}$, $\mathbf{Y}_{t}$ are sampled at specific ratios and then concatenated for updating. The sample indices are unified to ensure that points and their semantic labels correspond. To maintain $\mathbf{M}_{t}$ at a fixed size and prevent forgetting of early observations as input accumulates, we adjust the fusion ratios based on the timestep $t$. This ensures the total number of points in $\mathbf{M}_{t}$ remains below a predefined threshold, and the number of points sampled from each input frame is uniform.
@@ 4.3 Point Cloud Encoder and Semantic Encoder
Large-scale pre-training enables SpaitalLM to understand various indoor environments. However, we observed that SpatialLM’s performance degrades significantly when processing partially reconstructed scenes compared to full scene inputs. Therefore, we inject additional semantic information, predicted by pre-trained models, into the point cloud to aid the model in localizing and recognizing objects.
In our framework, the point cloud encoder follows the design of Sonata [51]. The point cloud $\mathbf{P}$ is encoded into 3D feature patches $\mathbf{F}$, where each patch represents a cubic region. To preserve the locality of point cloud features, our semantic encoder is designed with an identical pooling structure to the point cloud encoder, but it contains no trainable parameters between pooling operations, and only passes through a final linear projection layer for fine-tuning.
Specifically, for the semantic label of each point, we first convert it to its corresponding Llama-3.2-1B-Instruct [47] token feature. We then use the same pooling operations as the point cloud encoder to aggregate these per-point features into semantic feature patches. After that, we feed the patches into a linear projection layer, and add the resulting features to the point cloud features. This design not only ensures that the semantic information and point cloud features share the same granularity and spatial location, but also minimizes the number of trainable parameters, making fine-tuning possible even with limited 3D datasets.
@@ 5 Experiments
Method Avg. Fuzzy-$F_{1}$ Chair Table Computer Sink Bed Sofa Toilet
ScanNet++
SpatialLM-Ground-Truth-PCD 0.5708 0.7538 0.6160 0.3464 0.4078 0.8274 0.3883 0.5563
SpatialLM-Ground-Truth 0.6420 0.7541 0.6318 0.4540 0.6247 0.8263 0.4794 0.7233
SpatialLM-No-Finetune 0.0351 0.2126 0.0000 0.0194 0.0183 0.0823 0.0352 0.0559
SpatialLM-Merge 0.3397 0.5357 0.4606 0.1669 0.2320 0.4509 0.0273 0.3583
SpatialLM-Finetune 0.3943 0.5732 0.5160 0.2036 0.1451 0.4859 0.2897 0.2812
Ours 0.4397 0.6213 0.4717 0.2644 0.1683 0.6094 0.4633 0.4007
ScanNet
SpatialLM-Ground-Truth-PCD 0.5182 0.6050 0.5254 0.1464 0.4511 0.7482 0.6855 0.8803
SpatialLM-Ground-Truth 0.6138 0.6497 0.6123 0.3059 0.6063 0.9046 0.7113 0.9442
SpatialLM-No-Finetune 0.0383 0.1291 0.0000 0.0147 0.0470 0.1708 0.1089 0.1634
SpatialLM-Merge 0.2214 0.3267 0.2237 0.1125 0.1964 0.3517 0.3159 0.3065
SpatialLM-Finetune 0.2733 0.2900 0.2893 0.1322 0.2410 0.4743 0.3666 0.5651
Ours 0.3349 0.3425 0.3287 0.1621 0.2712 0.7166 0.4395 0.6198
Table 1: Quantitative Results of Our Experiments. We evaluate the methods on ScanNet [10] and ScanNet++ [58]. For each method, we report the average fuzzy $F_{1}$-score, as well as the fuzzy $F_{1}$-scores of 7 representative categories. Our method achieves state-of-the-art performance among the non-ground-truth baselines.
The primary goal of our experiments is to evaluate our system’s ability to continuously build and refine its understanding of novel environments from a monocular video stream. We aim to answer the following research questions: (1) How effectively does our online system perform in novel scenes? (2) Is maintaining a persistent spatial memory superior to simply merging independent per-frame predictions? (3) How does the integration of semantic information contribute to overall performance? (4) What is the performance impact of the external modules used in our system?
In this section, we first introduce the implementation details, baselines and datasets. We then present a comprehensive comparison, including both quantitative and qualitative results, to demonstrate the effectiveness of our method. Finally, we present ablation studies, to analyze some key design choices of our method.
@@ 5.1 Experimental Setup
Implementation Details In our system, we choose CUT3R [49] as our reconstruction module, and Grounded SAM [45] as our semantic annotator. For the point cloud encoder and the LLM backbone, we initialize them with SpatialLM1.1-Llama-1B [39]. During training, we only set the last layer of the point cloud encoder and the semantic encoder, as well as the LLM backbone as trainable, keeping all other parameters frozen. Training is conducted on 8 NVIDIA A800 GPUs, using standard cross-entropy loss for language model training.
@@ 5.2 Quantitative Results
As shown in Tab. 1, our method significantly outperforms the three lower-bound baselines on two benchmarks. The two upper-bound baselines, which use ground-truth data, establish the potential performance limit of our approach. This result demonstrates the overall effectiveness of our method. Furthermore, our system’s superior performance against SpatialLM-Merge highlights the importance of spatial memory maintaining, while the comparison with SpatialLM-Finetune confirms the benefits of injecting semantic information for object localization and recognition.

### v2

@@ Abstract
In recent years, researchers have increasingly been interested in how to enable Multimodal Large Language Models (MLLM) to possess spatial understanding and reasoning capabilities. However, most existing methods overlook the importance of the ability to continuously work in an ever-changing world, and lack the possibility of deployment on embodied systems in real-world environments. In this work, we introduce OnlineSI, a framework that can continuously improve its spatial understanding of its surroundings given a video stream. Our core idea is to maintain a finite spatial memory to retain past observations, ensuring the size of the spatial memory does not increase as the input accumulates. We further integrate 3D point cloud information with semantic information, helping MLLM to better locate and identify objects in the scene. To evaluate our method, we introduce the Fuzzy $F_{1}$-Score to mitigate ambiguity, and test our method on two representative datasets. Experiments demonstrate the effectiveness of our method, paving the way towards real-world embodied systems.
@@ 1 Introduction
However, equipping Multimodal Large Language Models (MLLM) with this form of online spatial intelligence remains challenging. First, most existing methods face prohibitive computational scaling as input stream grows. The architectures that repeatedly process all past observations with full attention layers quickly exhaust the model context and computational budget due to spatial and contextual redundancy. Second, many models still fail in understanding spatial relationships and reasoning in space, a limitation that is further exacerbated by the scarcity of 3D-grounded training data. While recent works, e.g., Hu et al. [hu20253dllm], have explored memory management for embodied planning, their solutions have two key limitations. As task execution proceeds, the memory bank infinitely grows, leading to computational and storage bottlenecks. Furthermore, their model perceives 3D space at a coarse-grained level, rendering them unsuitable for guiding fine-grained operations such as precise object manipulation. These difficulties motivate our core research question: Can multimodal large language models perceive and understand the 3D world in an online fashion?
In this paper, we propose OnlineSI, a framework designed to understand 3D scenes in an online fashion and make detections from streaming video. Our key insight is to maintain a finite, explicit spatial memory, and leverage the multimodal model to reason over this compact memory representation for precise, object-level predictions. This design ensures a maximum memory space limit and effectively curbs the growth of the computation budget. At the same time, the explicit spatial memory structure also facilitates the model’s grasp of fine-grained spatial concepts. Furthermore, we introduce a novel fusion technique that tightly integrates 3D point cloud data with semantic information, providing the MLLM with a rich, object-level understanding of the scene.
@@
Evaluating performance on an arbitrary video stream introduces a significant ambiguity: given the inherently incomplete observation, should a partially viewed object be counted as a detection? To address this evaluation challenge, we propose the Fuzzy $F_{1}$-Score, a modification of the standard $F_{1}$-Score designed for this online, partial observation setting. We conduct extensive experiments on two representative datasets, ScanNet [dai2017scannet] and ScanNet++ [yeshwanthliu2023scannetpp]. Experimental results show that our method significantly outperforms existing baselines, demonstrate its effectiveness in the online setting.
@@
  • We introduce OnlineSI, a novel framework for online 3D scene understanding and grounding that maintains bounded memory space and reduces the growth of inference cost, enabling it to process video streams incrementally.
@@ 2 Related Work
3D Object Detection aims to predict category labels and oriented 3D bounding boxes for each instance in the scene. Previous work can be divided into two main genres: methods that operate directly on 3D point clouds and those that use single-view 2D images. In the point-cloud-based genre, research has focused on enhancing the model’s understanding by enriching the 3D representation. A prominent strategy is to leverage knowledge from other modalities. For instance, [lu2022open, lu2023open, cao2023coda, cao2025collaborative] tap into knowledge priors from 2D images to achieve open-vocabulary 3D detection. Similarly, [peng2024global] employs a Large Language Model (LLM) to better process global-local information. Other approaches improve the model architecture or training strategy. For example, [shen2023v] uses a vertex relative positional encoding method to help the model focus on points near objects, while [kolodiazhnyi2025unidet3d] improves the system’s generalization through multi-dataset training. In the single-view-image-based genre, [brazil2023omni3d] pioneers unified monocular 3D object detection by training on the extensive Omni3D dataset. Following this, [li2025towards] introduces the first successful BEV-based monocular 3D object detector. A central theme in this area is addressing the scarcity of 3D training data by leveraging abundant 2D data. To this end, [huang2024training] uses open-vocabulary 2D models and pseudo-LiDAR to automatically label 3D objects. [yao2024open] leverages 2D results from Grounding DINO [liu2024grounding] with a 3D head to predict 3D bounding boxes. Both [jhang2025v] and [zhang2025detect] also address data scarcity by generating pseudo-3D data from 2D annotations or leveraging pre-trained 2D model knowledge. In our work, we emphasize on online detection from a video stream. A work with similar topic is [zhu2025move], which unifies grounding and exploration for embodied navigation. Compared to their work, we train a large language model to directly understand 3D point clouds and output 3D bounding boxes with rotations and class labels.
@@
Multimodal Large Language Models aim to understand, reason, and generate from multimodal information. Early methods, such as [radford2021learning], unifies images and texts by learning a joint image-text representation. More recently, with the remarkable success of Large Language Model and scaling laws, innumerable works have been devoted to unify vision and language within a single large transformer-based model [li2022blip, alayrac2022flamingo, li2023blip, liu2023visual, xie2024show, deitke2025molmo, bai2025qwen2, chen2025blip3, diao2025pixels]. Meanwhile, others have attempted to unify multiple modalities, enabling models to understand diverse information sources [hurst2024gpt, jiang2025solami, xu2025qwen3]. With the rapid development of autonomous driving and embodied AI, 3D understanding has recently become a significant focus for MLLM researchers. Some works aim to enable models to perceive spatial relationships from images by training on datasets containing 3D annotations [chen2024spatialvlm, cheng2024spatialrgpt, ma2025spatialllm, xu2025multi], while others seek to adapt models to 3D representations for spatial reasoning [fu2024scene, qi2024shapellm, xu2024pointllm, zhu2024llava, fan2025vlm, hu20253dllm, wu2025spatial, zheng2025learning, huang20253d]. However, most of the previous methods suffer from computational overload as the number of input images increases, making them unsuitable for deployment in embodied systems that need to learn at test time. Although some work has explored how to maintain a spatial memory for scene understanding [yang20253d, zou20253d], they do not focus on continuously improving understanding and online responding. In this sense, the work most similar to ours is [hu20253dllm]. However, they still have problems such as memory bank growing and coarse-level understanding. In our work, we update an explicit spatial memory based on input frames and use this memory as input to the MLLM, which guarantees a bounded memory space and fine-grained understanding.
@@
2D Object Detection is one of the most fundamental and challenging problems in computer vision. In earlier years, researchers typically use carefully designed, handcrafted features for detection [viola2001rapid, dalal2005histograms, felzenszwalb2008discriminatively]. Then with the rise of deep convolutional neural networks (CNN), many works introduced deep learning to this field, and they can be divided into two groups. The first one follows a “coarse-to-fine” process, proposing coarse regions and then focusing on them [girshick2014rich, he2015spatial, ren2015faster]. The second group, in contrast, retrieves all objects in one-step inference [redmon2016you, liu2016ssd, lin2017focal, law2018cornernet, carion2020end]. Recently, significant research has focused on open-vocabulary object detection, often achieved through large-scale pre-training [zang2022open, gu2021open, li2022grounded, liu2024grounding]. In our work, we use an off-the-shelf detector to obtain semantic labels. This semantic information is then injected into the point cloud to enhance its representation.
@@ 3 Preliminaries: SpatialLM
We leverage SpatialLM’s [mao2025spatiallm] ability of understanding point clouds to achieve online spatial intelligence. SpatialLM is a multimodal large language model designed to understand 3D indoor environment and generate structured scene descriptions. The model employs Sonata [wu2025sonata] as its point cloud encoder and finetunes Llama-3.2-1B-Instruct [touvron2023llama] on a synthetic dataset consisting of 12,328 distinct scenes.
@@ 4.1 Problem Formulation
Given a potentially infinite sequence of images $\{\mathbf{I}_{t}\in\mathbb{R}^{H\times W\times 3}\}_{t=1}^{N}$, for each current image $\mathbf{I}_{t}$, our goal is to predict the 3D bounding boxes and class labels for objects that have appeared in the scene, represented in a unified coordinate frame. We want the system to be capable of long-term learning from the environment; therefore, the computational and memory consumption of each inference step should not quickly become unbearable as the input accumulates. A good approach is to maintain a memory $\mathbf{M}_{t}$ of past observations, and ensure that $\mathbf{M}_{t}$ has a fixed upper bound on its size.
@@ 4.2 Spatial Memory Management
where $\mathbf{P}_{t-1}$, $\mathbf{X}^{\prime}_{t}$, $\mathbf{S}_{t-1}$, $\mathbf{Y}_{t}$ are sampled at specific ratios and then concatenated for updating. The sample indices are unified to ensure that points and their semantic labels correspond. To maintain $\mathbf{M}_{t}$ at a fixed size and prevent forgetting of early observations as input accumulates, we adjust the fusion ratios based on the timestep $t$. This ensures the total number of points in $\mathbf{M}_{t}$ remains below a predefined threshold, and the number of points sampled from each input frame is uniform. Regarding the effect of spatial memory management on reducing $\mathbf{M}_{t}$’s growth, please refer to Sec. 5.4.
@@ 4.3 Point Cloud Encoder and Semantic Encoder
Large-scale pre-training enables SpaitalLM to understand various indoor environments. However, we observed that SpatialLM’s performance degrades significantly when processing partially reconstructed scenes compared to full scene inputs. Therefore, we inject additional semantic information, which is predicted by pre-trained models, into the point cloud to aid the model in localizing and recognizing objects.
In our framework, the point cloud encoder follows the design of Sonata [wu2025sonata]. The point cloud $\mathbf{P}$ is encoded into 3D feature patches $\mathbf{F}$, where each patch represents a cubic region. To preserve the locality of point cloud features, our semantic encoder is designed with an identical pooling structure to the point cloud encoder, but it contains no trainable parameters between pooling operations, and only passes through a final linear projection layer for fine-tuning.
Specifically, for the semantic label of each point, we first convert it to its corresponding Llama-3.2-1B-Instruct [touvron2023llama] token feature. We then use the same pooling operations as the point cloud encoder to aggregate these per-point features into semantic feature patches. After that, we feed the patches into a linear projection layer, and add the resulting features to the point cloud features. This design not only ensures that the semantic information and point cloud features share the same granularity and spatial location, but also minimizes the number of trainable parameters, making fine-tuning possible with limited 3D datasets.
@@ 5 Experiments
The primary goal of our experiments is to evaluate our system’s ability to continuously build and refine its understanding of novel environments from a monocular video stream. We aim to answer the following research questions: (1) How effectively does our online system perform in novel scenes? (2) How well does our system perform in reducing the growth of inference computation and memory? (3) Is maintaining a persistent spatial memory superior to simply merging independent per-frame predictions? (4) How does the integration of semantic information contribute to overall performance? (5) What is the performance impact of the external modules used in our system?
In this section, we first introduce the implementation details, baselines and datasets. We then present a comprehensive comparison, including both quantitative and qualitative results, to demonstrate the effectiveness of our method. After that, we compare the inference cost scaling of our method with other spatial understanding benchmark, to illustrate how our system reduce the online inference cost. Finally, we present ablation studies, to analyze some key design choices of our method.
@@ 5.1 Experimental Setup
Implementation Details In our system, we choose CUT3R [wang2025continuous] as our reconstruction module, and Grounded SAM [ren2024grounded] as our semantic annotator. For the point cloud encoder and the LLM backbone, we initialize them with SpatialLM1.1-Llama-1B [mao2025spatiallm]. During training, we only set the last layer of the point cloud encoder and the semantic encoder, as well as the LLM backbone as trainable, keeping all other parameters frozen. Training is conducted on 8 NVIDIA A800 GPUs, using standard cross-entropy loss for language model training.
@@ 5.2 Quantitative Results
Method Avg. Fuzzy-$F_{1}$ Chair Table Computer Sink
[…]
SpatialLM-Ground-Truth-PCD 0.5708 0.7538 0.6160 0.3464 0.4078
SpatialLM-Ground-Truth 0.6420 0.7541 0.6318 0.4540 0.6247
SpatialLM-No-Finetune 0.0351 0.2126 0.0000 0.0194 0.0183
SpatialLM-Merge 0.3397 0.5357 0.4606 0.1669 0.2320
SpatialLM-Finetune 0.3943 0.5732 0.5160 0.2036 0.1451
[…]
SpatialLM-Ground-Truth-PCD 0.5182 0.6050 0.5254 0.1464 0.4511
SpatialLM-Ground-Truth 0.6138 0.6497 0.6123 0.3059 0.6063
SpatialLM-No-Finetune 0.0383 0.1291 0.0000 0.0147 0.0470
SpatialLM-Merge 0.2214 0.3267 0.2237 0.1125 0.1964
SpatialLM-Finetune 0.2733 0.2900 0.2893 0.1322 0.2410
[…]
Table 1: Quantitative Results of Our Experiments. We evaluate the methods on ScanNet [dai2017scannet] and ScanNet++ [yeshwanthliu2023scannetpp]. For each method, we report the average fuzzy $F_{1}$-score, as well as the fuzzy $F_{1}$-scores of 4 representative categories. Our method achieves state-of-the-art performance among the non-ground-truth baselines.
As shown in Tab. 5.2, our method significantly outperforms the three lower-bound baselines on two benchmarks. The two upper-bound baselines, which use ground-truth data, establish the potential performance limit of our approach. This result demonstrates the overall effectiveness of our method. Furthermore, our system’s superior performance against SpatialLM-Merge highlights the importance of spatial memory maintaining, while the comparison with SpatialLM-Finetune confirms the benefits of injecting semantic information for object localization and recognition.
5.3 Qualitative Results